tutor please help me solve for 5-(x-6)=4
Travis

Found 3 solutions by THANApHD, checkley71, scoop:
Answer by THANApHD(104)

5-(x-6)=4
5-x+6 =4
x = 4-11*-1
= -7*-1
=7

Answer by checkley71(8403)

5-(x-6)=4 or 5-x+6=4 or -x=4-5-6 or -x=-7 or x=7
proof 5-(7-6)=4 or 5-(1)=4 or 4=4

Answer by scoop(27)

5-(x-6)=4
5-x+6=4
5-x+6-6=4-6
5-x=-2
-5+5-x=2-5
-x=-3 divide both sides by negative (because x can not be neg.)
x=3/2